OpenStudy (anonymous):

Find the two products below. Compare and contrast, in complete sentences, the similarities and differences of the two. (x + 4)(x - 4) and (x + 4)(x + 4)

Nnesha (nnesha):

#first diff \[\huge\rm a^2 -b^2 = (a+b)(a-b)\] and \[\huge\rm (a+b)^2 = a^2 +2ab+b^2\]

Nnesha (nnesha):

a^2 -b^2 is a difference of square method which you can only apply if there is negative sign

OpenStudy (anonymous):

ok thanks

Nnesha (nnesha):

np